Flint Co. is building a new hockey arena at a cost of $2,660,000. It received a downpayment of $450,000 from local businessesDate Account Titles and Explanation Debit Credit January 1, 2019 Cash Bonds Payable Premium on Bonds Payable List of AccountsCash Paid Interest Expense Premium Amortization Carrying Amount of Bonds $ ta $ $Assume that on July 1, 2022, Flint Co. redeems half of the bonds at a cost of $1,172,400 plus accrued interest. Prepare the j
Answer #1
Face Amount $2,210,000
Interest Payment $2,210,000*11% =$243,100
Market Interest rate per period 10.00%
Cash Flow Table Value(PV of 10% for 10 years) Amount Present Value
PV of Interest 6.14457 $243,100 $1,493,745
PV of Principal 0.38554 $2,210,000 $852,043
PV of Bonds Payable(Issue Price) $2,345,788
Date Accounts and explanation Debit(in $) Credit(in $)
Jun 30,2017 Cash $                                    2,345,788
Bond Payable $                                     2,210,000
Premium on Bond Payable $                                        135,788
Col I Col II Col III Col IV
Date Cash Paid Interest expenses(Col V*3%) Premium amorrtization(Col I -Col II) Bond carrying amount
Jan 1,2019                               2,345,788
Jan 1,2020                                                   243,100                                           234,579                                                8,521                               2,337,267
Jan 1,2021                                                   243,100                                           233,727                                                9,373                               2,327,894
Jan 1,2022                                                   243,100                                           232,789                                              10,311                               2,317,583
Jan 1,2023                                                   121,550                                           115,880                                                5,670                               1,099,330
Interest expense accrued for 6 months period =$2,317,583*10%*6/12 =$115,880
Interest paid for 6 months =$121,550
Bond Premium amortized for 6 months =$121,550 - $115,880 =$5,670
Unamortized premium on Bond to be redeemed =$101,913/2 =$50,957
Date Accounts and explanation Debit(in $) Credit(in $)
Jul 1,2022 Interest expenses $                                       115,880
Premium on Bond Payable $                                           5,670
Cash $                                        121,550
July 1,2022 Bonds Payable $1,105,000
Premium on Bond payable $50,957
Loss on retirement of Bond $16,443
Cash $1,172,400